OpenStudy (anonymous):

2 + 2x = 8 solve for x?

OpenStudy (nincompoop):

I would start by isolating the term with x

OpenStudy (anonymous):

First, isolate x on one side of the equation, and then perform the arithmetic. \[2x = 8-2\] \[x = \frac{6}{2}\] \[x=3\]

OpenStudy (nincompoop):

2 -2 + 2x = 8 -2 2x = 6 then you want to find x not 2x so what to do next?